The series combination of five cells, each with electromotive force of 1.5 V and internal resistance of 0.5 Ω, is connected to an external resistance of 5.0 N. Find the (a) current through the external circuit, (b) terminal voltage of the battery, and (c) terminal voltage of a cell.


1
Expert's answer
2021-05-09T15:03:36-0400

Total Resistance of the circuit = R = R1 + r2 + R3 + R4 + R5 + R6

R = 0.5 + 0.5 + 0.5 + 0.5 + 0.5 + 5 = 7.5 Ohms


Total; EMF (or Voltage) V = E1 + E2 + E3 + E4 + E5

V = 1.5 + 1.5 + 1.5 + 1.5 + 1.5 = 7.5 V


Solution [a]

Current through the external circuit I "=\frac{V}{R}"

"I = \frac{7.5}{7.5}
= 1 Amp."


Solution [b]

As the five cells collectively make one battery. And all cells are in series, therefore,

Terminal voltage of the battery = E1 + E2 + E3 + E4 + E5

V = 1.5 + 1.5 + 1.5 + 1.5 + 1.5 = 7.5 V


Solution [c]

Terminal Voltage across individual/each Battery = 1.5V